OpenStudy (victoriaf16):

Write the equation of the line that passes through (1,5) and (-2,14) in slope-intercept form. Please help!!! thank you.

OpenStudy (anonymous):

do you know the formula that is used to create a straight line by using two points ?

OpenStudy (victoriaf16):

@esamalaa is it y2-y1/x2-x1?

OpenStudy (anonymous):

that is equal to? what is the other side of equation?

OpenStudy (anonymous):

slope-intercept form:\[y=mx+b\]where m is the slope and b is the y-intercept. you need to find the value of these or find another equation for the line that goes through your points and then solve that equation for y to get it into slope-intercept form.

OpenStudy (anonymous):

what you gave is the formula for finding the slope

OpenStudy (victoriaf16):

I'm so confused. I hate algebra.

OpenStudy (anonymous):

yea if she used this formula and then rearrange the equation to get the slope-intercept formula

OpenStudy (anonymous):

take a breath, relax and open your mind... i will help you. don't hate on algebra. a bad attitude never tkaes you anywhere good! you ready?

OpenStudy (victoriaf16):

Yes.

OpenStudy (anonymous):

 here are your points... connect them with a straight line

OpenStudy (victoriaf16):

unfortunately I can't even see your graph, half of it is cut out. :/

OpenStudy (anonymous):

point-slope form:\[y-y_1=m \left( x-x_1 \right)\]if you solve for y you get:\[y=mx-mx_1+y_1 \Rightarrow y=mx+\left( y_1-mx_1 \right) \Rightarrow \left(y_1-mx_1\right) = b\] \[m=\frac{ y_2-y_1 }{ x_2-x_1 }\]\[\left( x_1, y_1 \right) \text{ and } \left( x_2, y_2 \right) \text{ are points}\]